# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

### v2

@@ IV-B1 Problem Reformulation
Based on the objective function in (11a) and using the SINR expression in (2), we define
$\phi_{ik}({\bf e})\triangleq-\bigg{(}\frac{1}{\gamma_{ik}}\frac{|{\bf w}_{i}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}\bigg{)},$ (14)
for $k\in{\cal K}_{i}$, $i\in{\cal G}$. Furthermore, we denote the feasible set of $\widetilde{{\cal S}}_{{\bf w}}$ as ${\cal E}\triangleq\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$. Then, $\widetilde{{\cal S}}_{{\bf w}}$ can be rewritten as a min-max problem:
@@
The inner maximization problem is an integer program. To tackle this, let $\boldsymbol{\phi}({\bf e})\triangleq[\phi_{11}({\bf e}),\ldots,\phi_{1K_{1}}({\bf e}),\ldots,\phi_{G1}({\bf e}),\ldots,\phi_{GK_{G}}({\bf e})]^{T}$, and also let ${\bf y}$ be a $K_{\text{tot}}\times 1$ probability vector satisfying ${\bf y}\succcurlyeq 0$ and $\mathbf{1}^{T}{\bf y}=1$. Then, we can further transform problem (15) into the following equivalent problem
$\widetilde{{\cal S}}_{{\bf w}}^{\prime}:\min_{{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ f({\bf e},{\bf y})$ (16)
where $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$. Note that ${\cal Y}$ is a probability simplex. Thus, an optimal solution to the inner maximization problem in $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at some position that corresponds to $\max_{i,k}\phi_{ik}({\bf e})$, which is equivalent to the inner maximization in (15). Note that both ${\cal E}$ and ${\cal Y}$ are compact convex sets.
IV-B2 The Projected Subgradient Algorithm
Note that $f({\bf e},{\bf y})$ is concave in ${\bf y}$ but non-convex in ${\bf e}$. Thus, problem $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is a non-convex-concave min-max problem, which is typically NP-hard. Furthermore, the objective function $\max_{{\bf y}\in{\cal Y}}f({\bf e},{\bf y})$ may not be differentiable. To solve $\widetilde{{\cal S}}^{\prime}$, we notice that $\widetilde{{\cal S}}^{\prime}$ has the same structure as the min-max problem considered in [12] (i.e., with objective function $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and a convex feasible set). Thus, we adopt PSA in [12] to find a near-stationary point of $\widetilde{\cal S}^{\prime}_{{\bf w}}$.
Specifically, PSA is an iterative algorithm, with the updating procedure at iteration $j$ given by [^1] Note that PSA in [12] is provided using all real-valued variables. We provide the complex-valued version of this procedure, which can be shown to be equivalent to the real-valued version.
@@ IV-B2 The Projected Subgradient Algorithm
In (17), ${\bf y}^{(j)}$ can be obtained by finding index $\hat{i}\hat{k}$ such that $\phi_{\hat{i}\hat{k}}({\bf e})=\max_{i,k}\phi_{ik}({\bf e})$, and we have $f({\bf e},{\bf y}^{(j)})=\phi_{\hat{i}\hat{k}}({\bf e})$. To compute the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (18), with a slight abuse of notation to ease our presentation, we use $\phi_{ik}({\bf e})$ to represent $\phi_{i^{\prime}k^{\prime}}({\bf e})$ obtained in (17). Then, $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ has the following expression:
@@
$\quad=-\frac{{\bf q}_{\hat{i}\hat{i}\hat{k}}({\bf e})I_{\hat{i}\hat{k}}\!({\bf e})-\!\sum_{j\neq\hat{i}}{\bf q}_{j\hat{i}\hat{k}}({\bf e})S_{\hat{i}\hat{k}}({\bf e})}{\gamma_{\hat{i}\hat{k}}I_{\hat{i}\hat{k}}^{2}({\bf e})}-\frac{\zeta}{M}{\bf e}$ (20)
where $I_{\hat{i}\hat{k}}({\bf e})\triangleq\sum_{j\neq\hat{i}}|{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}+\sigma^{2}$, $S_{\hat{i}\hat{k}}({\bf e})\triangleq|{\bf w}_{\hat{i}}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}$, and ${\bf q}_{j\hat{i}\hat{k}}({\bf e})={\bf G}_{\hat{i}\hat{k}}^{H}{\bf w}_{j}{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})$.
The updates in each iteration are closed-form evaluations as in (17)–(IV-B2), with only matrix and vector multiplications. Thus, PSA is computationally cheap and is particularly suitable for solving $\widetilde{{\cal S}}_{{\bf w}}$ with a large value of $M$. Furthermore, the convergence analysis of PSA in [12] is applicable to our problem. It shows that the above PSA procedure is guaranteed to converge in finite time to a point at the vicinity of a stationary point of $\widetilde{{\cal S}}_{{\bf w}}$.
@@ IV-C Discussion on the AMBF Algorithm
We summarize our proposed fast AMBF algorithm for the RIS-assisted multicast QoS problem ${\cal P}_{o}$ in Algorithm 1. It consists of the proposed AMBF approach in Section III-B and the fast algorithms in Sections IV-A and IV-B for each subproblem. A few aspects of the algorithm are discussed below:
@@
1:Initialization: Set feasible initial point ${\bf e}^{0}$; Set $n=0$.
@@
3:     With ${\bf e}^{n}$, solve ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n+1}$ by using (13).
@@ IV-C1 Initialization
The AMBF algorithm requires an initial feasible point ${\bf e}^{0}$ that satisfies the constraints in (3c) for the BS multicast beamforming subproblem ${\cal P}_{\bf e}$. It can be easily obtained by generating a random phase for each element in ${\bf e}^{0}$. In each AO iteration $n$, solving $\tilde{{\cal S}}_{\bf w}$ by PSA also requires an initial point ${\bf e}^{(0)}$ for the update in (17). We should set ${\bf e}^{(0)}={\bf e}^{n}$, which is from the previous AO iteration and is also used in ${\cal P}_{\bf e}$ in the current AO iteration.
@@ IV-C2 Computational Complexity
Algorithm 1 consists of solving two optimization subproblems ${\cal P}_{{\bf e}}$ and $\widetilde{{\cal S}}_{{\bf w}}$ at each AO iteration. We discuss the computational complexity for solving each subproblem below:
Problem ${\cal P}_{\bf e}$
As mentioned in Section IV-A, using the optimal beamforming structure ${\bf w}$ in (13), ${\cal P}_{{\bf e}}$ is converted into the weight optimization problem w.r.t. $\{{\bf a}_{i}\}$ with total $K_{\text{tot}}(=\sum_{i=1}^{G}K_{i})$ variables, which can be solved by SCA [11]. If each SCA subproblem is solved using a typical interior point method [25] by the standard convex solver, the computational complexity is $O(K_{\text{tot}}^{3})$ per SCA iteration, which does not grow with either $M$ or $N$. We point out that this computational complexity can be further reduced by using the fast ADMM algorithm proposed recently in [26], which only involves closed-form updates and has the computational complexity in the order of $O(K_{\text{tot}}^{2})$.
Problem ${\cal S}_{{\bf w}}$
The problem is solved using PSA updates in (17) and (18) per iteration. Obtaining ${\bf y}^{(j)}$ in (17) involves computing $\textrm{SINR}_{ik}$ in (2) for all $K_{\text{tot}}$ users and finding the maximum $\phi_{ik}({\bf e})$, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+4)$ flops. The required computation in (18) is mainly at computing the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (IV-B2). We note that all the terms in (IV-B2), i.e., ${\bf q}_{j\hat{i}\hat{k}}$, $I_{\hat{i}\hat{k}}({\bf e})$, and $S_{\hat{i}\hat{k}}({\bf e})$ are part of the SINR expression that has already been computed. Thus, calculating $\nabla_{{\bf e}}\phi_{\hat{i}\hat{k}}({\bf e})$ only requires $2GM+2M$ flops. Thus, the entire updates per iteration requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+1)+2GM+2M$ flops, with the leading complexity being $K_{\text{tot}}N(G+M)$ flops. As typically we expect $M\gg G$, the leading computational complexity is $K_{\text{tot}}NM$. Thus, the complexity of PSA grows linearly over $M$, $N$, and $K_{\text{tot}}$. The complexity of the final per-element projection in (19) is $O(M)$. Thus, the overall computational complexity in each AO iteration of Algorithm 1 is as low as $O(K_{\text{tot}}^{2}+K_{\text{tot}}MN)$, which is linear in both $M$ and $N$.
@@ IV-C3 Convergence
V Generalization to RIS-Assisted Multi-user Downlink Beamforming
Problem ${\cal P}_{o}$ in fact represents the general formulation of downlink RIS-assisted multi-user beamforming QoS problem that includes both multicast and unicast transmission scenarios. Unicast transmission is the most commonly considered scenario in the downlink, where the BS sends data dedicated to each user. It is a special case of multicast beamforming when $K=1$ for each group, and the multicast beamforming problem is reduced to the conventional unicast beamforming for $G$ users. Thus, our proposed AMBF algorithm in Algorithm 1 can be directly used to solve the RIS-assisted multi-user unicast beamforming problem.
More importantly, we note that the transformed problem ${\cal P}_{2}$ in (III-A) reflects the general RIS-assisted downlink beamforming problem for either unicast or multicast transmissions. As a result, the AMBF approach in Section III-B to alternatingly solve the BS beamforming problem ${\cal P}_{\bf e}$ and the RIS passive beamforming problem ${\cal S}_{\bf w}$ is applicable in both scenarios.
So far, we have focused on the multicast beamforming problem ${\cal P}_{o}$ for $K>1$, which is a more challenging problem to solve than the unicast beamforming problem. For the unicast scenario, the RIS beamforming problem ${\cal S}_{\bf w}$ is still a multicast beamforming MMF problem as discussed in Remark 1 of Section III-B2, which can be efficiently solved by our proposed algorithm in Section IV-B. On the other hand, ${\cal P}_{\bf e}$ is reduced to the conventional BS multi-user beamforming problem w.r.t. ${\bf w}$ [27], which yields a closed-form optimal solution as briefly described below.
For $K=1$, we directly denote the single user in group $i\in{\cal G}$ as user $i$, with a total of $G$ users. The optimal beamforming solution ${\bf w}_{i}$ in (13) still holds as a special case, where the equivalent channel matrix $\widetilde{{\bf H}}_{i}$ for group $i$ is reduced to the equivalent channel $\tilde{{\bf h}}_{i}$ for user $i$, and weight vector ${\bf a}_{i}$ now becomes a scalar, denoted as $a_{i}$, i.e.,

### v3

@@ IV-B1 Problem Reformulation
We first transform $\widetilde{{\cal S}}_{\bf w}$ into a min-max problem formulation. Based on the objective function in (12a) and using the SINR expression in (1), we define
$\phi_{ik}({\bf e})\triangleq-\bigg{(}\frac{1}{\gamma_{ik}}\frac{|({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})^{H}{\bf w}_{j}|^{2}\!+\!\sigma^{2}}+\rho\frac{{\bf e}^{H}{\bf e}}{M}\bigg{)},$ (15)
for $k\in{\cal K}_{i}$, $i\in{\cal G}$. We denote the feasible set of $\widetilde{{\cal S}}_{{\bf w}}$ as ${\cal E}\triangleq\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$. Then, $\widetilde{{\cal S}}_{{\bf w}}$ can be rewritten as a min-max problem:
@@
To tackle the inner maximization problem, which is an integer program, let $\boldsymbol{\phi}({\bf e})$ be the $K_{\text{tot}}\times 1$ vector containing all $\phi_{ik}({\bf e})$’s, for $k\in{\cal K}_{i}$, $i\in{\cal G}$. Also let ${\bf y}$ be a $K_{\text{tot}}\times 1$ probability vector satisfying ${\bf y}\succcurlyeq 0$ and $\mathbf{1}^{T}{\bf y}=1$. Let $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$, and and ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$. Then, we can further transform problem (16) into the following equivalent problem
$\widetilde{{\cal S}}_{{\bf w}}^{\prime}:\min_{{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ f({\bf e},{\bf y}).$ (17)
Both ${\cal E}$ and ${\cal Y}$ are compact convex sets, and ${\cal Y}$ is a probability simplex. Thus, an optimal solution to the inner maximization problem in $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at a position that corresponds to $\max_{i,k}\phi_{ik}({\bf e})$, which is equivalent to the inner maximization in (16).
Problem $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is a nonconvex-concave min-max problem, since the objective function $f({\bf e},{\bf y})$ is concave in ${\bf y}$ but nonconvex in ${\bf e}$. Furthermore, function $g({\bf e})\triangleq\max_{{\bf y}\in{\cal Y}}f({\bf e},{\bf y})$ may not be differentiable w.r.t. ${\bf e}$, and thus, gradient-based methods are not applicable. We note that the structure of $\widetilde{{\cal S}}_{{\bf w}}^{\prime}$ is the same as the min-max problem structure considered in [21] (i.e., the objective function is in the form of $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y},$ and the feasible set is convex). It is shown in [21] that in this formulation, $f({\bf e},{\bf y})$ is an $L$-smooth function of ${\bf e}$, and $\nabla_{\bf e}f({\bf e},{\bf y})$ is a subgradient of $g({\bf e})$. In addition, $g({\bf e})$ is $L$-weakly convex over ${\cal E}$. Thus, we can adopt PSA [31] to find a near-stationary point of $\widetilde{\cal S}^{\prime}_{{\bf w}}$.
IV-B2 Projected Subgradient Algorithm
PSA is an iterative algorithm. Its updating procedure at iteration $j$ given by [^2] Note that PSA in [21] is provided using all real-valued variables. We provide the complex-valued version of this procedure, which can be shown to be equivalent.
@@ IV-B2 Projected Subgradient Algorithm
In (18), ${\bf y}^{(j)}$ can be obtained by finding an index pair $(\hat{i},\hat{k})$ such that $\phi_{\hat{i}\hat{k}}({\bf e})=\max_{i,k}\phi_{ik}({\bf e})$, and we have $f({\bf e},{\bf y}^{(j)})=\phi_{\hat{i}\hat{k}}({\bf e})$. The gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (19) has the following expression:
@@
$\quad=-\frac{{\bf q}_{\hat{i}\hat{i}\hat{k}}({\bf e})I_{\hat{i}\hat{k}}\!({\bf e})-\!\sum_{j\neq\hat{i}}{\bf q}_{j\hat{i}\hat{k}}({\bf e})S_{\hat{i}\hat{k}}({\bf e})}{\gamma_{\hat{i}\hat{k}}I_{\hat{i}\hat{k}}^{2}({\bf e})}-\frac{\rho}{M}{\bf e}$ (21)
where $I_{\hat{i}\hat{k}}({\bf e})\triangleq\sum_{j\neq\hat{i}}|({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})^{H}{\bf w}_{j}|^{2}+\sigma^{2}$, $S_{\hat{i}\hat{k}}({\bf e})\triangleq|({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})^{H}{\bf w}_{\hat{i}}|^{2}$, and ${\bf q}_{j\hat{i}\hat{k}}({\bf e})={\bf G}_{\hat{i}\hat{k}}^{H}{\bf w}_{j}{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})$.
The updates in each PSA iteration involve closed-form evaluations as in (18)–(IV-B2), with only matrix and vector multiplications. Thus, PSA is computationally cheap and is particularly suitable for solving $\widetilde{{\cal S}}_{{\bf w}}$ with a large value of $M$.
IV-B3 Convergence and Stopping Criterion
The convergence analysis of PSA in [21] is applicable to our problem. It shows that the above PSA procedure is guaranteed to converge in finite time to a point at the vicinity of a stationary point of $\widetilde{{\cal S}}_{{\bf w}}$.
Since PSA is a subgradient method, it is not guaranteed to move toward a descending direction. Thus, as commonly considered, we track the best solution obtained so far in $j$ iterations, i.e., the one with the minimum objective value of $\widetilde{{\cal S}}_{{\bf w}}^{\prime}$ in $j$ iterations: $g_{\min}^{(j)}=\min\{g^{(1)},...,g^{(j)}\}$, where $g^{(j)}=\max_{{\bf y}\in{\cal Y}}f({\bf e}^{(j)},{\bf y})$. Note that this is equivalent to track the maximum objective value in (12a) obtained so far. Based on its definition, $g_{\min}^{(j)}$ is monotonically decreasing and is bounded below. Thus, we can terminate the algorithm when $g_{\min}^{(j)}$ converges.
@@ IV-C Discussion on the AMBF Algorithm
We summarize our proposed fast AMBF algorithm for the RIS-assisted multicast QoS problem ${\cal P}_{o}$ in Algorithm 1. It consists of the proposed AMBF approach in Section III-B and the fast algorithms in Sections IV-A and IV-B for the two subproblems, respectively. A few aspects of the algorithm are discussed below:
@@
1:Initialization: Set feasible initial point ${\bf e}^{0}$; Set ${\rho}$ and $\alpha$; Set $n=0$.
@@
3:     With ${\bf e}^{n}$, solve ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n}$ by using (14).
@@ IV-C1 Initialization
The AMBF algorithm requires an initial feasible point ${\bf e}^{0}$ that satisfies the constraints in (2c) for the BS multicast beamforming subproblem ${\cal P}_{\bf e}$. It can be easily obtained by generating a random phase for each element in ${\bf e}^{0}$. In each AO iteration $n$, solving $\tilde{{\cal S}}_{\bf w}$ by PSA also requires an initial point ${\bf e}^{(0)}$ for the update in (18). We should set ${\bf e}^{(0)}={\bf e}^{n}$, i.e., the solution from the previous AO iteration. This is also used in ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n+1}$ in the current AO iteration.
@@ IV-C2 Computational Complexity
Algorithm 1 consists of solving two optimization subproblems ${\cal P}_{{\bf e}}$ and $\widetilde{{\cal S}}_{{\bf w}}$ at each iteration. The computational complexity for solving each subproblem is discussed below:
Problem ${\cal P}_{\bf e}$: As mentioned in Section IV-A, using the optimal beamforming structure ${\bf w}$ in (14), ${\cal P}_{{\bf e}}$ is converted into the weight optimization problem w.r.t. $\{{\bf a}_{i}\}$ with $K_{\text{tot}}(=\sum_{i=1}^{G}K_{i})$ variables, which can be solved by SCA [20]. If each SCA subproblem is solved using a typical interior point method [32] by the standard convex solver, the computational complexity is $O(K_{\text{tot}}^{3})$ per SCA iteration, which does not grow with either $M$ or $N$. We can adopt low-complexity fast ADMM-based algorithm [22], which only involves closed-form updates and has the computational complexity in the order of $O(K_{\text{tot}}^{2})$.
Problem ${\cal S}_{{\bf w}}$: The problem is solved using PSA updates in (18) and (19) per iteration. Obtaining ${\bf y}^{(j)}$ in (18) requires computing $\textrm{SINR}_{ik}$ in (1) for all $K_{\text{tot}}$ users and finding the maximum $\phi_{ik}({\bf e})$ in (15), which need $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+4)$ flops. The required computation in (19) is mainly at computing the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (IV-B2). We note that all the terms in (IV-B2), i.e., ${\bf q}_{j\hat{i}\hat{k}}$, $I_{\hat{i}\hat{k}}({\bf e})$, and $S_{\hat{i}\hat{k}}({\bf e})$ are part of the SINR expression that has already been computed. Thus, calculating $\nabla_{{\bf e}}\phi_{\hat{i}\hat{k}}({\bf e})$ only requires $2GM+2M$ flops. Thus, the entire updates per iteration requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+1)+2GM+2M$ flops, with the leading complexity being $K_{\text{tot}}N(G+M)$ flops. Since we expect $M\gg G$ in practical scenarios, the leading computational complexity is $K_{\text{tot}}NM$. Thus, the complexity of PSA for solving ${\cal S}_{\bf w}$ grows linearly over $M$, $N$, and $K_{\text{tot}}$.
The complexity of the final per-element projection in (20) is $O(M)$. The overall computational complexity in each AO iteration of Algorithm 1 is $O(K_{\text{tot}}^{2})+\text{const}\cdot K_{\text{tot}}MN$, which is linear in both $M$ and $N$ and quadratic in $K_{\text{tot}}$.
@@ IV-C3 Convergence
IV-D Special Case $K_{i}=1$
When $K_{i}=1$, $\forall i$, there is only one user in each group. Problem ${\cal P}_{o}$ is reduced to the RIS-assisted unicast scenario, where the BS transmits dedicated data to each of $G$ users
For this unicast scenario, the RIS passive beamforming problem ${\cal S}_{\bf w}$ is still a multicast beamforming MMF problem as in Section III-B2, which can be efficiently solved by the proposed PSA-based algorithm in Section IV-B2. The beamforming problem ${\cal P}_{\bf e}$ is reduced to the conventional downlink multi-user beamforming problem w.r.t. ${\bf w}$, for which a closed-form optimal solution can be obtained [33, 20]. In particular, SINR target for user $i$ is $\gamma_{i}$; for the optimal beamforming solution ${\bf w}_{i}$ in (14), the equivalent channel matrix $\widetilde{{\bf H}}_{i}$ for group $i$ reduces to an equivalent channel $\tilde{{\bf h}}_{i}$ for user $i$, and the weight vector ${\bf a}_{i}$ becomes a scalar, denoted as $a_{i}$. Then, we have